# Knowledge Augmentation in Federation: Rethinking What Collaborative Learning Can Bring Back to Decentralized Data

### v2

@@ IV-C Summary
In this section we revisit two representative lines of research that share the common idea of expanding user/client-specific domains of knowledge in a federation. Federated knowledge extrapolation mainly deals with data sparsity and missing information in local domains, whilst federated data augmentation focuses on mitigating non-IID local data distribution. The two approaches methodologically differ in the two aspects.
@@
  • The way new knowledge is introduced: knowledge extrapolation typically relies on representation learning to map multiple sources of information into one or several vector spaces, where sparse features are transformed into dense embeddings. These embeddings and the encoders act as the carriers in the process of knowledge aggregation and distribution. On this basis, predictions can be made to extend the local domain of knowledge, for example, by extrapolating unobserved interactions in the recommender system. By contrast, data augmentation can be performed independently to enhance the representations of knowledge, which is essentially analogous to linear interpolation [^7] In fact, it is exactly the idea behind the MixUp strategy.. Through collaboration, knowledge from outside the domain can be introduced by information fusion at sample level. This directly results in an expanded size of data and typically does not depend on any task-specific models.
  • Motivation and objectives: the extrapolation of knowledge is usually coupled with the optimization of task-specific objectives such as the accuracy of item recommendations. For example, in federated collaborative filtering, knowledge from different users are implicitly fused in the process of the global item vector update and local user vector update, guided by a cost function that quantifies the reconstruction loss. Federated data augmentation, on the other hand, aims to regulate the local data distribution towards the IID setting. In most cases, augmentation is performed without an explicit objective of optimization.
Each of the two approaches has certain limitations, as summarized in Table II. For collaborative knowledge extrapolation, domain-specific knowledge is embedded as factor matrices or vectors that are optimized under the supervision of known observations. This means that inaccurate information could be inferred and disseminated once the training samples are very noisy or manipulated by malicious clients. For collaborative data augmentation, neither interpolation-based or generative model-based schemes can theoretically guarantee the fidelity of synthetic samples. Certain mechanisms are necessary to constrain the process of augmentation against invalid patterns [60, 61]. So far, this problem still needs further investigation in the federated settings.
V Methods for Knowledge Filtering
A federated system is usually characterized by a rich set of data sources which, on the one hand, facilitate a collaborative way of learning but, on the other hand, bring about knowledge redundancy as well as the risk of data poisoning [62, 63]. The issue has become increasingly concerning with the advance of generative models that are able to fabricate adversarial samples. For example, a hyperdimensional computing (HDC) model is employed in [64] to craft poisoned samples that can bypass byzantine-robust defenses. This calls for more effective mechanisms for knowledge filtering.
We in this paper define knowledge filtering as the techniques for excluding misinformation in form of improper data (and its derivatives [^8] Including but not limited to embeddings, models and gradients.) and distinguishing different regions of the domain in terms of value for learning. Knowledge filtering can involve engineering and algorithm design at both levels across and within the domains. This section reviews the emerging studies that seek to address the discrepancy of knowledge regarding its validity and utility for the downstream task.
V-A Client-level Filtering
Poisoned datasets can severely threaten the validity of knowledge, which in turn causes performance degradation of the federated model learned. Being targeted or untargeted, data poisoning attacks are generally assumed to be carried out by one or multiple participants. Therefore, a straightforward approach to defending against data poisoning is to identify and exclude these malicious clients.
Although a variety of schemes have been developed for the generation of adversarial samples that are difficult to detect individually, they still inevitably change the distribution of data, for instance, with flipped labels or fake samples blended in. Through empirical studies on label-flipping attacks, Khuu et al. [65] found that, with a single round of attack, a small subset of adversarial clients can result in dramatic drop in the model’s recall on the targeted class as well as notable decrease of overall accuracy. Inspired by the Shapley Additive Explanation (SHAP), they developed a client filtering algorithm using Support Vector Machines (SVMs) to classify clients based on four features derived from SHAP values. However, SHAP values are computation-intensive for each client as we need to calculate a score matrix of all image-label pairs. By contrast, detecting data poisoning from the perspective of model is low-cost and more easy-to-implement. Model-based filtering approaches are based on the fact that local training on poisoned data will result in model update that are likely to deviate from that of benign clients. For example, the FedZZ [66] framework detects untargeted data poisoning attack based on the cosine similarity between a client’s update vector and a reference zone-aggregated vector. The proposed zone-based detection scheme groups the clients into multiple zones and calibrates the zone set progressively by accuracy. Similarly, Wei et al. [67] investigated the impact of targeted data poisoning attacks on local training by examining the eigenvalues of the covariances of gradient updates. Through theoretical analysis and empirical study, they proved that gradients from adversarial clients and honest clients form clearly separable clusters, which provides forensic evidence for accurate client filtering. Nonetheless, challenges remain as it was also observed that the timing of attacks greatly influences the effectiveness of statistic-based defense strategies [67].
Apart from adversarial clients with malicious intention, low-quality information such as noise and irrelevant samples may exist in local data. In our previous work [68], we conducted empirical studies inspecting the impact of low-quality information on the process of collaborative learning (Fig. 5). The result demonstrates that poor data conditions significantly slow down the convergence whilst undermining the global model’s performance.
@@ V-A Client-level Filtering
Figure 5: Results of our experiments over 100 clients under four disparate conditions of data that involve IID, non-IID and low-quality on-device data. Experiment details are described in [68], with more observations included.
@@
Instead of dirty labels, in [68] we examined the performance drop caused by low-quality clients that trained on local data affected by various forms of noise. We discovered that the engagement of low-quality knowledge can be as detrimental as pathological data distribution, and more importantly, the existence of both led to the worst results of convergence and final accuracy. To address the problem, a novel client filtering strategy FedProf [68] is presented, where the latent representations from the global model’s hidden layer are extracted and compressed into client-specific profiles. Based on the Kullback–Leibler (KL) divergence between the profiles, low-quality clients are filtered out before each FL round probabilistically.
@@
Noise data filtering at client level can be practically imperative especially for Internet of Things (IoT) applications. This is because real-time sensing is inherently susceptible to environmental disturbance and hardware flaws. For example, real-time data captured by image sensors are sensitive to violent camera shake and occlusion in the Internet of Vehicles (IoV) scenario. Targeting at this problem, Lei et al. [69] presented a filtering framework OES-Fed that works on the vehicle level. The framework first identifies a group of reference vehicles (i.e., vehicles with high-quality data) based on model accuracy, and sets their model parameters as the initial centers for K-means clustering. Vehicles that deviate significantly from the centers in the parameter space are detected as noise data carrier. In addition, Kalman filters and exponential smoothing are employed to incorporate the historical changes of model parameters. This enables dynamic assessment of local knowledge in response to Continual Learning (CL) scenarios [70] where new data keep rolling in.
V-B Sample-level Filtering
Studies on sample-level filtering endeavor to improve the quality of local knowledge by differentiating each sample within the local datasets. In contrast to the coarse-grained approaches that either keep or filter out entire local data domains, sample-level filtering corresponds to a fine-grained roadmap for enhancing the gain of local training and in turn facilitating global model convergence. However, we face unique challenges as it is typically not allowed in FL to access the raw attributes of each individual instance of data. This problem is especially intractable in the healthcare scenarios where the Electronic Health Records (EHRs) of patients are highly sensitive and probably heterogeneous across different organizations in the federation. By heterogeneous it means that samples from different clients scatter in inconsistent feature spaces that correspond to different permutations of features. To address this, Thakur et al. [71] proposed a knowledge abstraction and filtering framework, where the core component is the client-specific filtering modules that can be cascaded. Each filtering module starts with a native encoder that map local data to a unified latent space and ends with a gating operator that uses a globally shared knowledge abstraction vector to filter local samples’ representations. The novelty of this framework boils down to the knowledge abstraction vector which is trained collaboratively to embed consensus knowledge and guides the filtering of low knowledge. Nonetheless, the mechanism cannot directly identify and filter out problematic samples. Motivated by the knowledge distillation paradigm [72, 73], Lu et al. [74] proposed to employ conditional generators on each client to generate proxy samples for ensemble distillation. To address the loss of fidelity in synthetic data, the authors connected a filter to the head of each local model whose outputs are then modulated based on the reference labels. This helps in refining local knowledge without the need for a public data pool.
As a matter of fact, knowledge filtering can be performed simultaneously at client level and sample level. Adopting such a hybrid approach, Xu et al. [75] cast light on cloud-edge collaboration where “dirty” data on untrustworthy devices may significantly compromise the learning process. A novel framework called Safe is presented, which at the client level resorts to Alternating Direction Method of Multipliers (ADMM) for detecting suspicious clients that behave like data poisoning attackers. At sample level, data samples on these (potentially) malicious are then filtered based on K-means clustering and Euclidean distance. A shortcoming here is that the Euclidean distance in the raw feature space may not serve as a good criterion in the sample-level filtering in the presence of high dimensionality and sparsity. Moreover, Safe was only evaluated on the learning task of Support Vector machines (SVMs), which means that whether the framework can generalize to deep learning tasks remains questionable.
@@ V-B Sample-level Filtering
TABLE III: Summarizing the studies related to federated knowledge filtering in terms of methodological highlights, drawbacks, datasets for experiments, and the form of knowledge for exchange.
@@
[65] Client-level filtering derived features from SHAP $\bullet$ model-independency and statistical interpretability $\bullet$ high computational cost of SHAP MNIST and Fashion-MNIST
[66, 67] Client-level filtering gradients or update vectors $\bullet$ scalable and easy-to-implement
@@
[68] Client-level filtering statistically compressed representations $\bullet$ very low cost of communication
@@
[69] Client-level filtering task-specific models $\bullet$ scalable and effective against diverse types of data pollution
@@
[71] Sample-level filtering knowledge abstraction vector and the prediction model $\bullet$ abstraction of global knowledge for local filtering
@@
[75] Client+sample filtering model parameters $\bullet$ knowledge filtering at both levels using a multi-stage design
@@
[76] Fed. data re-weighting task-specific models $\bullet$ IWDS for achieving fast convergence without loss of accuracy $\bullet$ knowledge outside local domains unexploited CIFAR-10 and Fashion-MNIST
[77] Fed. data re-sampling task-specific models $\bullet$ easy to implement with any classic re-sampling methods $\bullet$ knowledge outside local domains unexploited MIMIC-III
[78] Fed. data re-weighting Gini coefficients and models $\bullet$ learnable strategy for client-specific re-weighting control
@@ V-C Federated Data Re-balancing
The notorious problem of non-IID data across the local domains in the federation not only hampers the convergence of the global model, but also leads to the drop of accuracy when the final model is evaluated on a presumably balanced test set. From the angle of knowledge, these phenomena can be explained from two aspects. First, theoretical evidence [79, 80] has revealed the convergence speed of the global model, on a certain ML task, is very much determined by the divergence of parameter updates (or accumulated gradients) from the clients. This essentially comes down to discrepancy in local knowledge as reflected by the conditional distribution of features and labels, which results in diverging local gradients towards different local optima. Second, non-IID data across the system usually implies that clients could have conflicting knowledge. Once learned and aggregated into a global model, client-specific knowledge gets diluted or neutralized, which consequently undermines the model’s ability to generalize.
Data re-balancing has proved its effectiveness in class-imbalanced learning where re-sampling is the most common technique adopted in the literature [81, 82]. By scaling up (down) the sampling probabilities of tail (dominant) classes, the local model is learned on a reshaped distribution of data. This can also be interpreted as a means of knowledge filtering where class-specific knowledge is kept by a different ratio. Considering that local data can be extremely imbalanced in non-IID settings, it is natural to explore collaborative knowledge filtering by means of data re-balancing. Tang et al. [76] investigated the impact of the celebrated class-balanced loss (CBLoss) [83] when applied to the local training of FL. Two interesting observations were made. First, data re-sampling narrowed the gap between the probability distributions of label sampling between different clients, which consequently accelerates the global model’s convergence. Second, though reaching the plateau at a faster rate, re-sampling resulted in an inferior model that yielded lower accuracy on local data. These two findings motivated them to devise a Imbalanced Weight Decay Sampling (IWDS) strategy that gradually smooths an imbalanced sampling towards uniform sampling. More specifically, given $n_{k,c}$ samples of class $c$ on client $u_{k}$, the sampling weight for a class-$c$ instance is:
@@
$\beta_{t}=\beta_{m}+(\beta_{0}-\beta_{m})\cdot\gamma^{t}$ (7)
@@
This strategy leverages imbalanced re-sampling to speed up the convergence at the early stage of FL and gradually approaches uniform sampling for higher final accuracy [76], i.e., $r^{(t)}_{k,i,j}$ gets closer to 1 with $\beta_{t}$ decreasing. Although distributional information is utilized locally to enhance local training, knowledge outside the local domain is left unexploited.
Ouyang et al. [78] formulated a different weighting rule for each sample as a power function of class frequency. After normalization the sampling probability $P_{k,c}$ for a class-$c$ instance is:
@@
where $C$ is total number of classes and the client-specific parameter $\beta_{k}$ controls the sampling probability distribution on client $u_{k}$. To dynamically determine $\{\beta_{k}|k=1,2,\ldots\}$ for the training cohort, a global coordinating agent based on Deep Reinforcement Learning (DRL) was developed. More specifically, the agent collects local and global model parameters as input state, makes decisions on $\{\beta_{k}|k=1,2,\ldots\}$, and receives rewards defined based on class-specific accuracy and Gini coefficients of local distribution.
Most of the classic re-sampling techniques for imbalanced learning can be easily applied to the local training step in FL. This corresponds to a naïve implementation of federated data re-balancing where clients act independently based on local distribution. Chen [77] conducted empirical studies to investigate the impacts of local data re-balancing on the efficacy of FL. Over-sampling, under-sampling and the combination of these two were evaluated on a healthcare-related logistic regression task over ten clients. They also filtered out the outliers in the pre-processing stage by z-score. Two interesting finds were reported: i) Under-sampling caused a slight performance decrease in the final model’s accuracy and precision, and ii) the hybrid method (combining over-sampling and under-sampling) showed the best performance across the board in all metrics. These observations imply a strong potential of data re-balancing in the federation even without any knowledge exchange between the domains. However, whether the phenomenon still holds for more complex tasks remains under-explored.

### v3

@@ IV-C Summary
Federated sample extrapolation mainly deals with data sparsity and missing information in local domains, whilst federated data augmentation focuses on mitigating non-IID local data distribution. We summarize related works in Table II and compare them in the following aspects:
@@
  • The way new knowledge is introduced: sample extrapolation typically relies on representation learning to map multiple sources of information into one or several vector spaces, using embeddings and encoder as the knowledge carriers. By contrast, data augmentation is essentially analogous to interpolation which fuses information at sample level. This directly results in an expanded size of data and typically does not depend on any task-specific models.
  • Motivation and objectives: the extrapolation of samples is sometimes coupled with the optimization of task-specific objectives such as the accuracy of item recommendations. Federated data augmentation, in many cases, can be performed in a task-agnostic fashion.
  • Limitations: For federated extrapolation, inaccurate information could be disseminated once the training samples are very noisy or manipulated by malicious clients. For augmentation, it is hard to theoretically guarantee the fidelity of synthetic samples.
V Collaborative Data Refinement
Federation can bring information redundancy as well as the risk of data poisoning to the participants [cina2024machine, Wang2023temporal]. We in this paper define data refinement as the techniques for excluding misinformation that exhibits knowledge incoherence.
@@ V Collaborative Data Refinement
Figure 5: (Left) results of our experiments over 100 clients under four disparate conditions of data that involve IID, non-IID and low-quality on-device data. (Right) the FedProf framework for dataset-level filtering. Experiment details are described in [wu2023fedprof], with more observations included.
V-A Dataset-level Filtering
A coarse-grained approach to filtering out misleading knowledge is to identify and exclude the malicious participants (and their datasets). Although tailored adversarial samples are difficult to detect individually, they still inevitably change the distribution of data. Through empirical studies on label-flipping attacks, Khuu et al. [Khuu2024shap] found that, with a single round of attack, a small subset of adversarial participants can result in dramatic drop in the model’s performance. Inspired by the Shapley Additive Explanation (SHAP), they developed a dataset filtering algorithm using Support Vector Machines (SVMs) to classify participants based on four features derived from SHAP values. However, SHAP values are computation-intensive for each participant as we need to calculate a score matrix of all image-label pairs. By contrast, Model-based filtering approaches are low-cost and more easy-to-implement, based on the fact that poisoned model updates are likely to deviate from benign updates. For example, the FedZZ [kumar2024precision] framework detects poisoned domains based on the cosine similarity between a client’s update vector and a reference zone-aggregated vector. The proposed zone-based detection scheme groups the participants into multiple zones and calibrates the zone set progressively by accuracy. Similarly, Wei et al. [Wei2024eigen] proposed to examine the eigenvalues of the covariance matrix of gradient updates. Through theoretical analysis and empirical study, they proved that gradients from adversarial participants and honest participants form clearly separable clusters, which provides forensic evidence for accurate dataset-level filtering. Li et al. [li2024mitigating] proposed a confidence-aware defense (CAD) strategy that is insensitive to the types of poisoning. By introducing the definition of confidence score and SuperLoss [castells2020superloss], the rationale behind CAD is that tampered data will cause a significant drop of confidence score in local training, based on which malicious participants can be detected.
@@ V-A Dataset-level Filtering
Apart from adversarial participants with malicious intention, low-quality information such as noise and irrelevant samples may exist in local data. In our previous work [wu2023fedprof], we empirically investigated the impact of low-quality information on the process of collaborative learning (left panel of Fig. 5). The result demonstrates that poor data conditions significantly slow down the convergence whilst undermining the global model’s performance. We discovered that the engagement of low-quality knowledge can be as detrimental as pathological data distribution. To address the problem, a novel participant filtering strategy FedProf [wu2023fedprof] is presented (right panel of Fig. 5). The key aspect of this strategy is termed “representation profiling” where the latent output from the global model’s hidden layer are extracted and compressed into participant-specific profiles. Based on the Kullback–Leibler (KL) divergence between the profiles, low-quality datasets are filtered out probabilistically before each FL round.
@@
Noise data filtering at dataset level can be practically imperative especially for Internet of Things (IoT) applications. This is because real-time sensing is inherently susceptible to environmental disturbance and hardware flaws. For example, real-time data captured by image sensors are sensitive to violent camera shake and occlusion in the Internet of Vehicles (IoV) scenario. Targeting at this problem, Lei et al. [lei2022oes] presented a filtering framework OES-Fed that works on the vehicle level. The framework first identifies a group of reference vehicles (i.e., vehicles with high-quality data) based on model accuracy, and sets their model parameters as the initial centers for K-means clustering. Vehicles that deviate significantly from the centers in the parameter space are detected as noise data carrier. In addition, Kalman filters and exponential smoothing are employed to incorporate the historical changes of model parameters. This enables dynamic assessment of local knowledge in response to Continual Learning (CL) scenarios [liang2025diffusion] where new data keep rolling in.
@@
TABLE III: Summarizing the studies related to collaborative data refinement in terms of methodological highlights, drawbacks, datasets for experiments, and the form of knowledge for exchange.
@@
[Khuu2024shap] Dataset-level filtering derived features from SHAP $\bullet$ model-independency and statistical interpretability $\bullet$ high computational cost of SHAP MNIST and Fashion-MNIST
[kumar2024precision, Wei2024eigen] Dataset-level filtering gradients or update vectors $\bullet$ scalable and easy-to-implement
@@
[wu2023fedprof] Dataset-level filtering statistically compressed representations $\bullet$ very low cost of communication
@@
[lei2022oes] Dataset-level filtering task-specific models $\bullet$ scalable and effective against diverse types of data pollution
@@
[thakur2024knowledge] Sample-level filtering knowledge abstraction vector and the prediction model $\bullet$ abstraction of global knowledge for local filtering
@@
[xu2022safe] Dataset+sample filtering model parameters $\bullet$ knowledge filtering at both levels using a multi-stage design
@@
[tang2021data] Federated data re-weighting task-specific models $\bullet$ IWDS for achieving fast convergence without loss of accuracy $\bullet$ knowledge outside local domains unexploited CIFAR-10 and Fashion-MNIST
[chen2024implementing] Federated data re-sampling task-specific models $\bullet$ easy to implement with any classic re-sampling methods $\bullet$ knowledge outside local domains unexploited MIMIC-III
[ouyang2024ddpg] Federated data re-weighting Gini coefficients and models $\bullet$ learnable strategy for client-specific re-weighting control
@@
V-B Sample-level Filtering
In contrast to the coarse-grained approaches that either keep or filter out entire local data domains, sample-level filtering identifies the deviants in each domain.
However, it is typically not allowed in a federated setting to access the raw attributes of each individual data instances. This problem is especially intractable in the healthcare scenarios where the Electronic Health Records (EHRs) of patients are highly sensitive and probably heterogeneous across different organizations in the federation. By “heterogeneous” it means that samples from different clients scatter in inconsistent feature spaces that correspond to different permutations of features. To address this, Thakur et al. [thakur2024knowledge] proposed a knowledge abstraction and filtering framework, where the core component is the participant-specific filtering modules that can be cascaded. Each filtering module starts with a native encoder that map local data to a unified latent space and ends with a gating operator that uses a globally shared knowledge abstraction vector to filter local samples’ representations. Nonetheless, the mechanism cannot directly identify and filter out problematic samples. Motivated by the knowledge distillation paradigm [le2025fedmekt, chen2023metafed], Lu et al. [lu2024datafree] proposed to employ local conditional generators to generate proxy samples for ensemble distillation. A filter is connected to the head of each local model whose outputs are then modulated based on the reference labels. This helps in refining local knowledge without the need for a public data pool.
As a matter of fact, data refinement can be performed simultaneously at dataset level and sample level. Adopting such a hybrid approach, Xu et al. [xu2022safe] cast light on cloud-edge collaboration where “dirty” data on untrustworthy devices may significantly compromise the learning process. A novel framework called Safe is presented, which at the dataset level resorts to Alternating Direction Method of Multipliers (ADMM) for detecting suspicious participants that behave like data poisoning attackers. At sample level, data samples on these (potentially) malicious are then filtered based on K-means clustering and Euclidean distance. A shortcoming here is that the Euclidean distance in the raw feature space may not serve as a good criterion due to high dimensionality and sparsity. Moreover, Safe was only evaluated on the learning task of Support Vector machines (SVMs), which means that whether the framework can generalize to deep learning tasks remains questionable.
@@ V-C Federated Data Re-balancing
Data imbalance, usually associated with lopsided marginal probabilities $P(X)$ and $P(Y)$, creates difficulties for knowledge discovering. Data re-balancing has proved its effectiveness in class-imbalanced learning where re-sampling is the most common technique adopted in the literature [shi2023advances, zhou2025heterogeneous]. By scaling up (down) the sampling probabilities of tail (dominant) classes, the local model is learned on a reshaped distribution of data. This can also be interpreted as a means of knowledge filtering where class-specific knowledge is kept by a different ratio. Tang et al. [tang2021data] investigated the impact of the celebrated class-balanced loss (CBLoss) [cui2019cbloss] when applied to the federated setting. Two interesting observations were made. First, data re-sampling narrowed the gap of the probability distributions of label sampling between different clients. Second, though reaching the plateau at a faster rate, re-sampling can result in an inferior model that yielded lower accuracy on local data. These two findings motivated them to devise a Imbalanced Weight Decay Sampling (IWDS) strategy that gradually smooths an imbalanced sampling towards uniform sampling. More specifically, given $n_{k,c}$ samples of class $c$ on client $u_{k}$, the sampling weight for a class-$c$ instance is:
@@
$\beta_{t}=\beta_{m}+(\beta_{0}-\beta_{m})\cdot\gamma^{t},$ (5)
@@
This strategy leverages imbalanced re-sampling to speed up the convergence at the early stage of collaboration and gradually approaches uniform sampling for higher final accuracy [tang2021data], i.e., $r^{(t)}_{k,i,j}$ gets closer to 1 with $\beta_{t}$ decreasing. Although distributional information is utilized locally to enhance local training, knowledge outside the local domain is left unexploited.
Ouyang et al. [ouyang2024ddpg] formulated a different weighting rule for each sample as a power function of class frequency. After normalization the sampling probability $P_{k,c}$ for a class-$c$ instance is:
@@
where $C$ is total number of classes and the participant-specific parameter $\beta_{k}$ controls the sampling probability distribution on participant $u_{k}$. To dynamically determine $\{\beta_{k}|k=1,2,\ldots\}$ for the training cohort, a global coordinating agent based on Deep Reinforcement Learning (DRL) was developed. More specifically, the agent collects local and global model parameters as input state, makes decisions on $\{\beta_{k}|k=1,2,\ldots\}$, and receives rewards defined based on class-specific accuracy and Gini coefficients of local distribution.
Most of the classic re-sampling techniques for imbalanced learning can be easily applied to the local operation phase in TAF. This corresponds to a naïve implementation of federated data re-balancing where clients act independently based on local distribution. Chen [chen2024implementing] conducted empirical studies to investigate its impacts on some downstream learning tasks. Over-sampling, under-sampling and the combination of these two were evaluated. Two interesting finds were reported: i) Under-sampling caused a slight performance loss in the final model’s accuracy and precision, and ii) the hybrid method (combining over-sampling and under-sampling) showed the best performance across the board in all metrics. These observations imply a strong potential of data re-balancing in the federation even without any knowledge exchange between the domains. However, whether the phenomenon still holds for more complex tasks remains under-explored.
We summarize these solutions in Table III.